# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ Appendix C Results of varying the input nuclear reactions
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.22$ $-0.06$ ${}^{168}\mathrm{Yb}$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.22$ $-0.10$
${}^{124}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.20$ $-0.07$ ${}^{166}\mathrm{Yb}(\gamma,n)$ $-0.18$ $-0.07$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.36$ $-0.18$ ${}^{166}\mathrm{Yb}(\gamma,\alpha)$ $-0.34$ $-0.12$
${}^{125}\mathrm{Xe}(\gamma,n)$ $-0.54$ $-0.33$ ${}^{167}\mathrm{Yb}(\gamma,n)$ $0.19$ $0.07$
${}^{126}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.19$ $-0.07$ ${}^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.43$ $0.16$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.39$ $-0.18$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.29$ $0.10$
@@
${}^{127}\mathrm{Xe}(\gamma,n)$ $-0.57$ $-0.29$ ${}^{174}\mathrm{Hf}$ ${}^{172}\mathrm{Hf}(\gamma,\alpha)$ $-0.31$ $-0.06$
${}^{130}\mathrm{Ba}$ ${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.16$ $-0.02$ ${}^{174}\mathrm{W}(\gamma,n)$ $-0.19$ $-0.03$
${}^{131}\mathrm{Ba}(\gamma,n)$ $-0.81$ $-0.15$ ${}^{174}\mathrm{W}(\gamma,\alpha)$ $-0.30$ $-0.07$
@@
${}^{133}\mathrm{Ba}(\gamma,n)$ $-0.74$ $-0.27$ ${}^{182}\mathrm{Os}(\gamma,\alpha)$ $0.17$ $0.03$
${}^{138}\mathrm{La}$ ${}^{133}\mathrm{La}(\gamma,p)$ $-0.19$ $-0.08$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.04$
${}^{135}\mathrm{La}(\gamma,n)$ $-0.31$ $-0.12$ ${}^{180}\mathrm{Ta}$ ${}^{179}\mathrm{Ta}(\gamma,n)$ $-0.89$ $-0.07$
${}^{136}\mathrm{La}(\gamma,n)$ $0.25$ $0.08$ ${}^{180}\mathrm{W}$ ${}^{180}\mathrm{Os}(\gamma,n)$ $-0.26$ $-0.08$
${}^{137}\mathrm{La}(\gamma,n)$ $-0.35$ $-0.16$ ${}^{180}\mathrm{Os}(\gamma,\alpha)$ $-0.50$ $-0.21$
${}^{136}\mathrm{Ce}$ ${}^{137}\mathrm{Ce}(\gamma,n)$ $-0.47$ $-0.07$ ${}^{181}\mathrm{Os}(\gamma,n)$ $-0.16$ $-0.04$
${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.31$ $-0.04$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.22$ $0.06$
${}^{138}\mathrm{Nd}(\gamma,p)$ $0.52$ $0.07$ ${}^{184}\mathrm{Os}$ ${}^{182}\mathrm{Os}(\gamma,\alpha)$ $-0.19$ $-0.02$
@@
${}^{138}\mathrm{Ce}$ ${}^{137}\mathrm{Ce}(\gamma,n)$ $0.66$ $0.08$ ${}^{184}\mathrm{Pt}(\gamma,\alpha)$ $-0.45$ $-0.04$
${}^{139}\mathrm{Ce}(\gamma,n)$ $-0.37$ $-0.04$ ${}^{186}\mathrm{Pt}(\gamma,\alpha)$ $-0.16$ $-0.01$
${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.16$ $-0.02$ ${}^{188}\mathrm{Pt}(\gamma,\alpha)$ $0.16$ $0.01$
@@
${}^{140}\mathrm{Nd}(\gamma,\alpha)$ $0.18$ $0.02$ ${}^{190}\mathrm{Pt}$ ${}^{190}\mathrm{Hg}(\gamma,n)$ $-0.30$ $-0.04$
${}^{190}\mathrm{Hg}(\gamma,\alpha)$ $-0.51$ $-0.07$
@@
${}^{196}\mathrm{Hg}$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $-0.68$ $-0.42$
${}^{197}\mathrm{Pb}(\gamma,n)$ $-0.28$ $-0.13$
@@
Table C4: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the 10×10$\times$10 ×3D-inspired mixing scenario.
@@
${}^{74}\mathrm{Se}$ ${}^{75}\mathrm{Se}(\gamma,n)$ $-0.87$ $-0.30$ ${}^{138}\mathrm{Ce}$ ${}^{137}\mathrm{Ce}(\gamma,n)$ $0.58$ $0.05$
${}^{78}\mathrm{Kr}$ ${}^{79}\mathrm{Kr}(\gamma,n)$ $-0.81$ $-0.35$ ${}^{139}\mathrm{Ce}(\gamma,n)$ $-0.19$ $-0.01$
${}^{84}\mathrm{Sr}$ ${}^{84}\mathrm{Rb}(\gamma,n)$ $0.16$ $0.06$ ${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.31$ $-0.02$
${}^{85}\mathrm{Sr}(\gamma,n)$ $-0.81$ $-0.33$ ${}^{138}\mathrm{Nd}(\gamma,p)$ $0.37$ $0.03$
${}^{92}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $-0.93$ $-0.17$ ${}^{138}\mathrm{Nd}(\gamma,\alpha)$ $-0.17$ $-0.01$
@@
${}^{96}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $-0.88$ $-0.25$ ${}^{144}\mathrm{Sm}$ ${}^{142}\mathrm{Sm}(\gamma,n)$ $-0.31$ $-0.04$
${}^{98}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $0.68$ $0.07$ ${}^{142}\mathrm{Sm}(\gamma,p)$ $-0.29$ $-0.03$
${}^{100}\mathrm{Pd}(\gamma,p)$ $0.45$ $0.05$ ${}^{143}\mathrm{Sm}(\gamma,n)$ $-0.41$ $-0.05$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.29$ $-0.03$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.35$ $0.04$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.18$ $0.02$ ${}^{152}\mathrm{Gd}$ ${}^{152}\mathrm{Dy}(\gamma,\alpha)$ $-0.50$ $-0.02$
${}^{102}\mathrm{Pd}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $-0.29$ $-0.08$ ${}^{154}\mathrm{Dy}(\gamma,\alpha)$ $-0.16$ $-0.00$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.30$ $-0.08$ ${}^{158}\mathrm{Er}(\gamma,n)$ $0.15$ $0.00$
${}^{103}\mathrm{Pd}(\gamma,n)$ $-0.65$ $-0.21$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.19$ $0.00$
${}^{106}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $-0.83$ $-0.30$ ${}^{156}\mathrm{Dy}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.19$ $-0.06$
@@
${}^{109}\mathrm{Cd}(\gamma,n)$ $-0.50$ $-0.10$ ${}^{158}\mathrm{Dy}$ ${}^{159}\mathrm{Dy}(\gamma,n)$ $0.19$ $0.01$
${}^{113}\mathrm{In}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.17$ $-0.05$ ${}^{158}\mathrm{Er}(\gamma,n)$ $-0.16$ $-0.01$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.24$ $-0.08$ ${}^{158}\mathrm{Er}(\gamma,\alpha)$ $-0.28$ $-0.01$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.82$ $0.34$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.18$ $-0.01$
${}^{112}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.17$ $-0.05$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.56$ $0.02$
${}^{113}\mathrm{Sn}(\gamma,n)$ $-0.78$ $-0.33$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.01$
${}^{114}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.18$ $-0.02$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.01$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.27$ $-0.04$ ${}^{162}\mathrm{Er}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.28$ $-0.10$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.76$ $0.12$ ${}^{159}\mathrm{Er}(\gamma,\alpha)$ $-0.16$ $-0.05$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.15$ $0.03$ ${}^{160}\mathrm{Er}(\gamma,n)$ $-0.19$ $-0.07$
${}^{169}\mathrm{Lu}(\gamma,n)$ $0.16$ $0.02$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $-0.29$ $-0.09$
${}^{115}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.18$ $-0.02$ ${}^{161}\mathrm{Er}(\gamma,n)$ $0.18$ $0.06$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.27$ $-0.04$ ${}^{166}\mathrm{Yb}(\gamma,\alpha)$ $0.40$ $0.12$
@@
${}^{120}\mathrm{Te}$ ${}^{119}\mathrm{Te}(\gamma,n)$ $-0.20$ $-0.04$ ${}^{164}\mathrm{Er}$ ${}^{164}\mathrm{Yb}(\gamma,n)$ $-0.23$ $-0.10$
${}^{121}\mathrm{Te}(\gamma,n)$ $-0.66$ $-0.13$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.59$ $-0.43$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.42$ $0.07$ ${}^{165}\mathrm{Yb}(\gamma,n)$ $-0.17$ $-0.09$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.35$ $-0.06$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.07$
${}^{124}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.22$ $-0.08$ ${}^{168}\mathrm{Yb}$ ${}^{168}\mathrm{Hf}(\gamma,n)$ $-0.16$ $-0.05$
${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.27$ $-0.10$ ${}^{168}\mathrm{Hf}(\gamma,\alpha)$ $-0.40$ $-0.21$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.44$ $-0.23$ ${}^{169}\mathrm{Hf}(\gamma,n)$ $-0.17$ $-0.05$
${}^{125}\mathrm{Xe}(\gamma,n)$ $-0.40$ $-0.22$ ${}^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.41$ $0.19$
${}^{126}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,n)$ $-0.21$ $-0.06$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.22$ $0.08$
${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.26$ $-0.08$ ${}^{174}\mathrm{Hf}$ ${}^{174}\mathrm{W}(\gamma,n)$ $-0.23$ $-0.06$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.46$ $-0.19$ ${}^{174}\mathrm{W}(\gamma,\alpha)$ $-0.52$ $-0.19$
${}^{125}\mathrm{Xe}(\gamma,n)$ $0.38$ $0.16$ ${}^{176}\mathrm{W}(\gamma,\alpha)$ $-0.15$ $-0.05$
${}^{127}\mathrm{Xe}(\gamma,n)$ $-0.26$ $-0.11$ ${}^{180}\mathrm{Ta}$ ${}^{179}\mathrm{Ta}(\gamma,n)$ $-0.88$ $-0.03$
${}^{130}\mathrm{Ba}$ ${}^{126}\mathrm{Ba}(\gamma,p)$ $-0.18$ $-0.02$ ${}^{180}\mathrm{W}$ ${}^{180}\mathrm{Os}(\gamma,n)$ $-0.30$ $-0.12$
${}^{126}\mathrm{Ba}(\gamma,\alpha)$ $-0.31$ $-0.03$ ${}^{180}\mathrm{Os}(\gamma,\alpha)$ $-0.54$ $-0.39$
${}^{128}\mathrm{Ba}(\gamma,n)$ $-0.26$ $-0.03$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.21$ $0.08$
${}^{128}\mathrm{Ba}(\gamma,p)$ $-0.16$ $-0.01$ ${}^{184}\mathrm{Os}$ ${}^{184}\mathrm{Pt}(\gamma,\alpha)$ $-0.49$ $-0.10$
${}^{128}\mathrm{Ba}(\gamma,\alpha)$ $-0.27$ $-0.02$ ${}^{185}\mathrm{Pt}(\gamma,\alpha)$ $-0.16$ $-0.02$
@@
${}^{131}\mathrm{Ba}(\gamma,n)$ $-0.51$ $-0.05$ ${}^{190}\mathrm{Pt}$ ${}^{190}\mathrm{Hg}(\gamma,n)$ $-0.27$ $-0.06$
${}^{132}\mathrm{Ba}$ ${}^{128}\mathrm{Ba}(\gamma,\alpha)$ $-0.18$ $-0.02$ ${}^{190}\mathrm{Hg}(\gamma,\alpha)$ $-0.52$ $-0.15$
@@
${}^{133}\mathrm{Ba}(\gamma,n)$ $-0.56$ $-0.10$ ${}^{196}\mathrm{Hg}$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $-0.75$ $-0.53$
${}^{138}\mathrm{La}$ ${}^{137}\mathrm{La}(\gamma,n)$ $-0.65$ $-0.35$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.12$
${}^{136}\mathrm{Ce}$ ${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.39$ $-0.06$
@@
${}^{138}\mathrm{Nd}(\gamma,\alpha)$ $-0.16$ $-0.03$
@@
Table C5: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the 50×50$\times$50 ×3D-inspired mixing scenario.

### v2

@@ Appendix C Results of varying the input nuclear reactions
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.22-0.22 0.06-0.06 $^{168}\mathrm{Yb}$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.22-0.22 0.10-0.10
$^{124}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,p)$ 0.20-0.20 0.07-0.07 $^{166}\mathrm{Yb}(\gamma,n)$ 0.18-0.18 0.07-0.07
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.36-0.36 0.18-0.18 $^{166}\mathrm{Yb}(\gamma,\alpha)$ 0.34-0.34 0.12-0.12
$^{125}\mathrm{Xe}(\gamma,n)$ 0.54-0.54 0.33-0.33 $^{167}\mathrm{Yb}(\gamma,n)$ $0.19$ $0.07$
$^{126}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,p)$ 0.19-0.19 0.07-0.07 $^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.43$ $0.16$
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.39-0.39 0.18-0.18 $^{196}\mathrm{Pb}(\gamma,n)$ $0.29$ $0.10$
@@
$^{127}\mathrm{Xe}(\gamma,n)$ 0.57-0.57 0.29-0.29 $^{174}\mathrm{Hf}$ $^{172}\mathrm{Hf}(\gamma,\alpha)$ 0.31-0.31 0.06-0.06
$^{130}\mathrm{Ba}$ $^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.16-0.16 0.02-0.02 $^{174}\mathrm{W}(\gamma,n)$ 0.19-0.19 0.03-0.03
$^{131}\mathrm{Ba}(\gamma,n)$ 0.81-0.81 0.15-0.15 $^{174}\mathrm{W}(\gamma,\alpha)$ 0.30-0.30 0.07-0.07
@@
$^{133}\mathrm{Ba}(\gamma,n)$ 0.74-0.74 0.27-0.27 $^{182}\mathrm{Os}(\gamma,\alpha)$ $0.17$ $0.03$
$^{138}\mathrm{La}$ $^{133}\mathrm{La}(\gamma,p)$ 0.19-0.19 0.08-0.08 $^{196}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.04$
$^{135}\mathrm{La}(\gamma,n)$ 0.31-0.31 0.12-0.12 $^{180}\mathrm{Ta}$ $^{179}\mathrm{Ta}(\gamma,n)$ 0.89-0.89 0.07-0.07
$^{136}\mathrm{La}(\gamma,n)$ $0.25$ $0.08$ $^{180}\mathrm{W}$ $^{180}\mathrm{Os}(\gamma,n)$ 0.26-0.26 0.08-0.08
$^{137}\mathrm{La}(\gamma,n)$ 0.35-0.35 0.16-0.16 $^{180}\mathrm{Os}(\gamma,\alpha)$ 0.50-0.50 0.21-0.21
$^{136}\mathrm{Ce}$ $^{137}\mathrm{Ce}(\gamma,n)$ 0.47-0.47 0.07-0.07 $^{181}\mathrm{Os}(\gamma,n)$ 0.16-0.16 0.04-0.04
$^{138}\mathrm{Nd}(\gamma,n)$ 0.31-0.31 0.04-0.04 $^{196}\mathrm{Pb}(\gamma,n)$ $0.22$ $0.06$
$^{138}\mathrm{Nd}(\gamma,p)$ $0.52$ $0.07$ $^{184}\mathrm{Os}$ $^{182}\mathrm{Os}(\gamma,\alpha)$ 0.19-0.19 0.02-0.02
@@
$^{138}\mathrm{Ce}$ $^{137}\mathrm{Ce}(\gamma,n)$ $0.66$ $0.08$ $^{184}\mathrm{Pt}(\gamma,\alpha)$ 0.45-0.45 0.04-0.04
$^{139}\mathrm{Ce}(\gamma,n)$ 0.37-0.37 0.04-0.04 $^{186}\mathrm{Pt}(\gamma,\alpha)$ 0.16-0.16 0.01-0.01
$^{138}\mathrm{Nd}(\gamma,n)$ 0.16-0.16 0.02-0.02 $^{188}\mathrm{Pt}(\gamma,\alpha)$ $0.16$ $0.01$
@@
$^{140}\mathrm{Nd}(\gamma,\alpha)$ $0.18$ $0.02$ $^{190}\mathrm{Pt}$ $^{190}\mathrm{Hg}(\gamma,n)$ 0.30-0.30 0.04-0.04
$^{190}\mathrm{Hg}(\gamma,\alpha)$ 0.51-0.51 0.07-0.07
@@
$^{196}\mathrm{Hg}$ $^{196}\mathrm{Pb}(\gamma,n)$ 0.68-0.68 0.42-0.42
$^{197}\mathrm{Pb}(\gamma,n)$ 0.28-0.28 0.13-0.13
@@
Table C4: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the $10\times$3D-inspired mixing scenario.
@@
$^{74}\mathrm{Se}$ $^{75}\mathrm{Se}(\gamma,n)$ 0.87-0.87 0.30-0.30 $^{138}\mathrm{Ce}$ $^{137}\mathrm{Ce}(\gamma,n)$ $0.58$ $0.05$
$^{78}\mathrm{Kr}$ $^{79}\mathrm{Kr}(\gamma,n)$ 0.81-0.81 0.35-0.35 $^{139}\mathrm{Ce}(\gamma,n)$ 0.19-0.19 0.01-0.01
$^{84}\mathrm{Sr}$ $^{84}\mathrm{Rb}(\gamma,n)$ $0.16$ $0.06$ $^{138}\mathrm{Nd}(\gamma,n)$ 0.31-0.31 0.02-0.02
$^{85}\mathrm{Sr}(\gamma,n)$ 0.81-0.81 0.33-0.33 $^{138}\mathrm{Nd}(\gamma,p)$ $0.37$ $0.03$
$^{92}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ 0.93-0.93 0.17-0.17 $^{138}\mathrm{Nd}(\gamma,\alpha)$ 0.17-0.17 0.01-0.01
@@
$^{96}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ 0.88-0.88 0.25-0.25 $^{144}\mathrm{Sm}$ $^{142}\mathrm{Sm}(\gamma,n)$ 0.31-0.31 0.04-0.04
$^{98}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ $0.68$ $0.07$ $^{142}\mathrm{Sm}(\gamma,p)$ 0.29-0.29 0.03-0.03
$^{100}\mathrm{Pd}(\gamma,p)$ $0.45$ $0.05$ $^{143}\mathrm{Sm}(\gamma,n)$ 0.41-0.41 0.05-0.05
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.29-0.29 0.03-0.03 $^{196}\mathrm{Pb}(\gamma,n)$ $0.35$ $0.04$
$^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.18$ $0.02$ $^{152}\mathrm{Gd}$ $^{152}\mathrm{Dy}(\gamma,\alpha)$ 0.50-0.50 0.02-0.02
$^{102}\mathrm{Pd}$ $^{100}\mathrm{Pd}(\gamma,p)$ 0.29-0.29 0.08-0.08 $^{154}\mathrm{Dy}(\gamma,\alpha)$ 0.16-0.16 0.00-0.00
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.30-0.30 0.08-0.08 $^{158}\mathrm{Er}(\gamma,n)$ $0.15$ $0.00$
$^{103}\mathrm{Pd}(\gamma,n)$ 0.65-0.65 0.21-0.21 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.19$ $0.00$
$^{106}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ 0.83-0.83 0.30-0.30 $^{156}\mathrm{Dy}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.19-0.19 0.06-0.06
@@
$^{109}\mathrm{Cd}(\gamma,n)$ 0.50-0.50 0.10-0.10 $^{158}\mathrm{Dy}$ $^{159}\mathrm{Dy}(\gamma,n)$ $0.19$ $0.01$
$^{113}\mathrm{In}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.17-0.17 0.05-0.05 $^{158}\mathrm{Er}(\gamma,n)$ 0.16-0.16 0.01-0.01
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.24-0.24 0.08-0.08 $^{158}\mathrm{Er}(\gamma,\alpha)$ 0.28-0.28 0.01-0.01
$^{113}\mathrm{Sn}(\gamma,n)$ $0.82$ $0.34$ $^{159}\mathrm{Er}(\gamma,n)$ 0.18-0.18 0.01-0.01
$^{112}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.17-0.17 0.05-0.05 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.56$ $0.02$
$^{113}\mathrm{Sn}(\gamma,n)$ 0.78-0.78 0.33-0.33 $^{196}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.01$
$^{114}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.18-0.18 0.02-0.02 $^{202}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.01$
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.27-0.27 0.04-0.04 $^{162}\mathrm{Er}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.28-0.28 0.10-0.10
$^{113}\mathrm{Sn}(\gamma,n)$ $0.76$ $0.12$ $^{159}\mathrm{Er}(\gamma,\alpha)$ 0.16-0.16 0.05-0.05
$^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.15$ $0.03$ $^{160}\mathrm{Er}(\gamma,n)$ 0.19-0.19 0.07-0.07
$^{169}\mathrm{Lu}(\gamma,n)$ $0.16$ $0.02$ $^{160}\mathrm{Er}(\gamma,\alpha)$ 0.29-0.29 0.09-0.09
$^{115}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.18-0.18 0.02-0.02 $^{161}\mathrm{Er}(\gamma,n)$ $0.18$ $0.06$
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.27-0.27 0.04-0.04 $^{166}\mathrm{Yb}(\gamma,\alpha)$ $0.40$ $0.12$
@@
$^{120}\mathrm{Te}$ $^{119}\mathrm{Te}(\gamma,n)$ 0.20-0.20 0.04-0.04 $^{164}\mathrm{Er}$ $^{164}\mathrm{Yb}(\gamma,n)$ 0.23-0.23 0.10-0.10
$^{121}\mathrm{Te}(\gamma,n)$ 0.66-0.66 0.13-0.13 $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.59-0.59 0.43-0.43
$^{122}\mathrm{Xe}(\gamma,p)$ $0.42$ $0.07$ $^{165}\mathrm{Yb}(\gamma,n)$ 0.17-0.17 0.09-0.09
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.35-0.35 0.06-0.06 $^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.07$
$^{124}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,n)$ 0.22-0.22 0.08-0.08 $^{168}\mathrm{Yb}$ $^{168}\mathrm{Hf}(\gamma,n)$ 0.16-0.16 0.05-0.05
$^{122}\mathrm{Xe}(\gamma,p)$ 0.27-0.27 0.10-0.10 $^{168}\mathrm{Hf}(\gamma,\alpha)$ 0.40-0.40 0.21-0.21
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.44-0.44 0.23-0.23 $^{169}\mathrm{Hf}(\gamma,n)$ 0.17-0.17 0.05-0.05
$^{125}\mathrm{Xe}(\gamma,n)$ 0.40-0.40 0.22-0.22 $^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.41$ $0.19$
$^{126}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,n)$ 0.21-0.21 0.06-0.06 $^{196}\mathrm{Pb}(\gamma,n)$ $0.22$ $0.08$
$^{122}\mathrm{Xe}(\gamma,p)$ 0.26-0.26 0.08-0.08 $^{174}\mathrm{Hf}$ $^{174}\mathrm{W}(\gamma,n)$ 0.23-0.23 0.06-0.06
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.46-0.46 0.19-0.19 $^{174}\mathrm{W}(\gamma,\alpha)$ 0.52-0.52 0.19-0.19
$^{125}\mathrm{Xe}(\gamma,n)$ $0.38$ $0.16$ $^{176}\mathrm{W}(\gamma,\alpha)$ 0.15-0.15 0.05-0.05
$^{127}\mathrm{Xe}(\gamma,n)$ 0.26-0.26 0.11-0.11 $^{180}\mathrm{Ta}$ $^{179}\mathrm{Ta}(\gamma,n)$ 0.88-0.88 0.03-0.03
$^{130}\mathrm{Ba}$ $^{126}\mathrm{Ba}(\gamma,p)$ 0.18-0.18 0.02-0.02 $^{180}\mathrm{W}$ $^{180}\mathrm{Os}(\gamma,n)$ 0.30-0.30 0.12-0.12
$^{126}\mathrm{Ba}(\gamma,\alpha)$ 0.31-0.31 0.03-0.03 $^{180}\mathrm{Os}(\gamma,\alpha)$ 0.54-0.54 0.39-0.39
$^{128}\mathrm{Ba}(\gamma,n)$ 0.26-0.26 0.03-0.03 $^{196}\mathrm{Pb}(\gamma,n)$ $0.21$ $0.08$
$^{128}\mathrm{Ba}(\gamma,p)$ 0.16-0.16 0.01-0.01 $^{184}\mathrm{Os}$ $^{184}\mathrm{Pt}(\gamma,\alpha)$ 0.49-0.49 0.10-0.10
$^{128}\mathrm{Ba}(\gamma,\alpha)$ 0.27-0.27 0.02-0.02 $^{185}\mathrm{Pt}(\gamma,\alpha)$ 0.16-0.16 0.02-0.02
@@
$^{131}\mathrm{Ba}(\gamma,n)$ 0.51-0.51 0.05-0.05 $^{190}\mathrm{Pt}$ $^{190}\mathrm{Hg}(\gamma,n)$ 0.27-0.27 0.06-0.06
$^{132}\mathrm{Ba}$ $^{128}\mathrm{Ba}(\gamma,\alpha)$ 0.18-0.18 0.02-0.02 $^{190}\mathrm{Hg}(\gamma,\alpha)$ 0.52-0.52 0.15-0.15
@@
$^{133}\mathrm{Ba}(\gamma,n)$ 0.56-0.56 0.10-0.10 $^{196}\mathrm{Hg}$ $^{196}\mathrm{Pb}(\gamma,n)$ 0.75-0.75 0.53-0.53
$^{138}\mathrm{La}$ $^{137}\mathrm{La}(\gamma,n)$ 0.65-0.65 0.35-0.35 $^{202}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.12$
$^{136}\mathrm{Ce}$ $^{138}\mathrm{Nd}(\gamma,n)$ 0.39-0.39 0.06-0.06
@@
$^{138}\mathrm{Nd}(\gamma,\alpha)$ 0.16-0.16 0.03-0.03
@@
Table C5: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the $50\times$3D-inspired mixing scenario.